# Design of Bayesian Clinical Trials with Clustered Data and Multiple Endpoints

### v2

@@ 6 Discussion
In this paper, we proposed an efficient framework to assess design operating characteristics for Bayesian clinical trials with clustered data. This framework determines the minimum number of clusters required to ensure the probability of making a correct decision is sufficiently large while bounding the probability of making an incorrect decision. The computational efficiency of our framework is motivated by considering a proxy for the joint sampling distribution of posterior probabilities across multiple estimands. We use the behaviour in this large-sample proxy distribution to justify estimating true sampling distributions at only two sample sizes. Our method therefore drastically reduces the number of simulation repetitions required to design Bayesian trials. Furthermore, we repurpose our estimates of the sampling distribution to construct bootstrap confidence intervals that quantify the impact of simulation variability on the sample size recommendation. Our methodology can be broadly used in clinical settings where large-sample regularity conditions are satisfied – including settings with a single estimand or those that do not require marginalization via Bayesian G-computation.
Our motivating example in Section 5 provides insight for future research. First, our example reflects that clinicians often want to consider the trial operating characteristics for various data generation processes $\Psi_{1}$ and $\Psi_{0}$. We already efficiently explore the sample size space, but analogs to Theorem 1 that enable efficient consideration of the $\Psi_{1}$-space and $\Psi_{0}$-space for a given sample size would be useful. Moreover, we analyzed the data in our simulated trials using separate models for each binary outcome. This choice was made for simplicity, but it would be worthwhile to develop Bayesian models for analyzing dependent binary outcomes that do not compromise the computational feasibility of trial design. When multiple outcomes inform complex decision criteria in clinical trials, it is also important to propose more sophisticated procedures to choose decision thresholds than the ad-hoc methods used in this paper. These extensions could help control design operating characteristics in settings where some but not all of the subhypotheses $\{H_{0,k}\}_{k=1}^{K}$ are true. We also provided a preliminary extension of our methods to accommodate adaptive designs, but this extension could be formalized and further developed in future work.
Lastly, the use of dynamic borrowing methods to incorporate information from prior trials has become increasingly common. Depending on the borrowing method used, the conditions for the BvM theorem may not be satisfied. Further research is needed to clarify when these conditions are met and develop extensions for scenarios where the conditions are not satisfied.
@@
These materials include proofs of Lemma 1 and Theorem 1, as well as additional context for the example in Section 5. The code to conduct the numerical studies in the paper is available online: https://github.com/lmhagar/ClusterDOCs.
@@ Acknowledgements
The authors would like to thank Dr. Dick Menzies for his comments and suggestions regarding the SSTaRLeT application component of this article.

### v3

@@ 8 Discussion
In this paper, we proposed an efficient framework to assess design operating characteristics for Bayesian clinical trials with clustered data. This framework determines the minimum number of clusters required to ensure power (or assurance) is sufficiently large while bounding the expected type I error rate. The computational efficiency of our framework is motivated by considering a proxy for the sampling distribution of posterior probabilities. We use the behaviour in this large-sample proxy distribution to justify estimating true sampling distributions at only two cluster counts. Our method, therefore, drastically reduces the number of simulation repetitions required to design Bayesian trials. Furthermore, we repurpose our estimates of the sampling distribution to construct bootstrap confidence intervals that quantify the impact of simulation variability on the cluster count recommendation. Our methodology can be broadly used in clinical settings where large-sample regularity conditions are satisfied – including settings that do not require marginalization via Bayesian G-computation.
Our design methodology efficiently recommends the number of clusters $c$ for trials with clustered data. However, we acknowledge that the cluster count is only one aspect of the total sample size in cluster trials. We assign a (potentially degenerate) probability distribution for the cluster sizes $\{n_{j}\}_{j=1}^{c}$ by specifying the auxiliary data-generation process $\boldsymbol{\zeta}$. Our recommendations for $c$ thus depend on these choices. To consider sample size recommendations with different cluster sizes, we must reimplement our algorithm. Additional simulations would also be required to consider different cluster sizes in a standard simulation-based design framework. Furthermore, we emphasize that the asymptotic results that substantiate our methods apply when the number of clusters $c\rightarrow\infty$. This theory does not apply when there are a small number of clusters with large cluster sizes.
The SSTARLET trial that inspired our illustrative example provides insights for future research described in Section 7. Moreover, our example reflects that clinicians often want to consider the trial operating characteristics when parameters $\boldsymbol{\theta}$ are generated according to various probability models $\Psi_{1}$ and $\Psi_{0}$. We already efficiently explore the sample size space, but analogs to Theorem 1 that enable efficient consideration of the $\Psi_{1}$-space and $\Psi_{0}$-space for a given cluster count would be useful. Lastly, the use of dynamic borrowing methods to incorporate information from prior trials has become increasingly common. Depending on the borrowing method used, the conditions for the BvM theorem may not be satisfied. Further research is needed to clarify when these conditions are met and develop extensions for scenarios where the conditions are not satisfied.
@@
These materials include proofs of Lemma 1 and Theorem 1, as well as additional context for the example in Section 6 and additional numerical studies. The code to conduct the numerical studies in the paper is available online: https://github.com/lmhagar/ClusterDOCs.
@@ Acknowledgements
The authors would like to thank Dr. Dick Menzies for his comments regarding the SSTARLET trial that inspired the illustrative example for this paper.